# Anomalous Coherence Length in Superconductors with Quantum Metric

### v5

@@ Abstract
The coherence length $\xi$ is the fundamental length scale of superconductors which governs the sizes of Cooper pairs, vortices, Andreev bound states, and more. In existing microscopic theories of superconductivity, it is expected that as the strength of the attractive interaction increases, $\xi$ decreases as the electrons are bound together more strongly. In BCS theory, the coherence length is $\xi_{\mathrm{BCS}}=\hbar v_{F}/\Delta$, where $v_{F}$ is the Fermi velocity and $\Delta$ is the pairing gap. It is clear that increasing $\Delta$ will shorten $\xi_{\mathrm{BCS}}$. However, the situation is puzzling for superconductors with exactly flat bands in which $v_{F}$ goes to zero and $\xi_{\mathrm{BCS}}$ is expected to be zero. In this work, we show that the quantum metric, which is the real part of the quantum geometric tensor, gives rise to an anomalous contribution to the coherence length. Specifically, $\xi=\sqrt{\xi_{\mathrm{BCS}}^{2}+\ell_{\mathrm{qm}}^{2}}$ for a superconductor where $\ell_{\mathrm{qm}}$ is the quantum metric contribution. In the flat-band limit, $\xi$ does not vanish but is bound below by $\ell_{\mathrm{qm}}$. We demonstrate that under the uniform pairing condition, $\ell_{\mathrm{qm}}$ is controlled by the quantum metric of minimal trace in the flat-band limit. Physically, the Cooper pair size of a superconductor cannot be squeezed down to a size smaller than $\ell_{\mathrm{qm}}$ which is a fundamental length scale determined by the quantum geometry of the wave functions. Lastly, we compute the quantum metric contributions for the family of superconducting moiré graphene materials, demonstrating the significant role played by quantum metric effects in these narrow-band superconductors.
@@ I Introduction
Bardeen-Cooper-Schriffer (BCS) theory Bardeen et al. (1957) of superconductivity stands as one of the most successful and influential theories in modern physics. It offers a mean-field, yet non-perturbative and microscopic framework for understanding superconductivity. It has been very successful in describing a large number of superconductors Carbotte (1990); Sigrist and Ueda (1991); Blatter et al. (1994). Deviations from the BCS theory are not unusual, which are often attributed to strong interaction effects Stewart (1984); Keimer et al. (2015). Recently, the observations of superconductivity in twisted bilayer graphene Cao et al. (2018); Yankowitz et al. (2019); Arora et al. (2020); Oh et al. (2021); Tian et al. (2023) and related graphene family Liu et al. (2020); Park et al. (2021, 2022) hinted that a new theory is needed to describe superconductors with nearly flat bands. It was observed in a recent experiment that some important physical quantities deviate greatly from BCS predictions and the microscopic origins behind them are not yet clear. For example, the BCS superconducting coherence length $\xi_{\mathrm{BCS}}$ is expressed as $\hbar v_{F}/\Delta$, where $v_{F}$ is the Fermi velocity and $\Delta$ is the pairing gap. When the moiré band of twisted bilayer graphene is nearly flat with $v_{F}\approx 10^{3}\text{m}/\text{s}$ and $\Delta\approx 0.2$ meV, $\xi_{\mathrm{BCS}}$ is estimated to be around $3$ nm which is more than one order of magnitude shorter than the values measured using upper critical field measurements Tian et al. (2023). Furthermore, the low Fermi velocity (or equivalently, large effective mass) should lead to a low superfluid stiffness. This results in an expected Berezinskii-Kosterlitz-Thouless transition temperature much lower than the transition temperature measured at optimal doping Tian et al. (2023). It had been pointed out by previous works that the quantum metric Hu et al. (2019); Julku et al. (2020) of the flat bands, which is the real part of the quantum geometric tensor, is essential in sustaining a supercurrent Peotta and Törmä (2015); Törmä et al. (2022). Apart from the investigations of superfluid weight Julku et al. (2016); Liang et al. (2017a, b); Iskin (2018a, b, c); Mondaini et al. (2018); Iskin (2019); Hu et al. (2019); Julku et al. (2020); Xie et al. (2020); Verma et al. (2021); Herzog-Arbeitman et al. (2022a); Kitamura et al. (2022); Huhtinen et al. (2022a); Mao and Chowdhury (2024); Hofmann et al. (2023a); Mao and Chowdhury (2023), the quantum metric affects other physical quantities such as the intrinsic nonlinear transport Gao et al. (2023); Wang et al. (2023); Kaplan et al. (2024) and electron-phonon coupling Yu et al. (2024).
@@
Explicitly, considering the Bloch states of a band represented by $|u(\bm{k})\rangle$, the quantum geometric tensor $\mathfrak{G}$  Provost and Vallee (1980a); Berry (1984) is
@@
Here, $a$ and $b$ represent the momentum directions. The quantum geometric tensor can be decomposed into real and imaginary parts as $\mathfrak{G}=\mathcal{G}-i\mathcal{F}/2$, where the real part $\mathcal{G}$ is the quantum metric and the imaginary part $\mathcal{F}$ is the Berry curvature. Berry curvature arises from the phase difference between adjacent Bloch states and characterizes the band topology of materials Berry (1984); Klitzing et al. (1980); Thouless et al. (1982); Bellissard et al. (1994); Hasan and Kane (2010); Qi and Zhang (2011); Berry (1989); Bouhon et al. (2023). The study of the physical consequences of the Berry curvature has been one of the central topics in modern physics. On the other hand, the effect of quantum metric, which measures the distance between two quantum states Anandan and Aharonov (1990); Resta (2011), is much less studied. It was pointed out that the quantum metric provides the size (or the so-called quadratic spread) $\ell_{\mathrm{qm}}$ of the optimally localized Wannier state of a band Marzari and Vanderbilt (1997), where $\ell_{\mathrm{qm}}=\sqrt[4]{\det\overline{\mathcal{G}}}$. Here, $\overline{\mathcal{G}}$, defined in Eq. (11), is the weighted average of the quantum metric of the Bloch states within a band. The mathematical definition of $\ell_{\mathrm{qm}}$, which we call the quantum metric length, measures the minimal spread of the Wannier functions, is schematically illustrated in Fig. 1. However, the impact of the quantum metric length $\ell_{\mathrm{qm}}$ on physical quantities was not clear. Until very recently, the Ginzburg-Landau theory Chen and Law (2024) shown that at zero temperature, $\xi=\ell_{\mathrm{qm}}$ for an exactly flat band, which is independent of the interaction strength.
@@
In other words, there is an anomalous quantum metric contribution to the superconducting coherence length (recall that $\xi_{\mathrm{BCS}}=\hbar v_{F}/\Delta$). In the flat-band limit with vanishing $v_{F}$, the quantum metric effect can be significant and even dominant. We show that this is indeed the case for several moiré superconductors with nearly flat bands Tian et al. (2023); Park et al. (2021, 2022); Liu et al. (2020). Our result gives a possible explanation for why the observed superconducting coherence length in the recent experiment Tian et al. (2023) is much larger than expected. It is worth noting that the coherence length is lattice-geometry independent while the quantum metric is lattice-geometry dependent Simon and Rudner (2020); Huhtinen et al. (2022b). To resolve the discrepancy, we apply the uniform pairing condition when evaluating the pair correlators and then demonstrate that $\ell_{\mathrm{qm}}$ is related to the quantum metric of the minimal trace Huhtinen et al. (2022b). We delineate the physical picture that, in the presence of the quantum metric, increasing the attractive interaction strength between electrons can only reduce the BCS part of the coherence length and squeeze the Cooper pair size down to the quantum metric length $\ell_{\mathrm{qm}}$, but not further, as demonstrated in Fig. 2.
Figure 2: a For a conventional superconductor with a dispersive band (as illustrated by the insert) without quantum metric, the coherence length $\xi=\hbar v_{F}/\Delta$ decreases as $\Delta$ increases and $\xi$ is not bounded from below. b In the presence of quantum metric, the superconducting coherence length $\xi$ has a lower bound of $\ell_{\mathrm{qm}}$. For a superconductor with a narrow band (as illustrated in the insert), the conventional contribution can be suppressed as $\Delta$ increases.
Additionally, for a topological flat band with nontrivial (spin) Chern number, $\xi\geq a\sqrt{|C|/4\pi}$, where $a$ is the lattice constant as illustrated in Fig. 4. At the end of this work, we show that the quantum metric length $\ell_{\mathrm{qm}}$ is important for the superconducting moiré graphene family. As the new length scale $\ell_{\mathrm{qm}}$ defined by the quantum metric is a fundamental property of the band structure and its importance should be manifested beyond superconducting phenomena, we expect that $\ell_{\mathrm{qm}}$ also plays a crucial role in other interaction-driven ordered states (such as the magnetic or density-wave states Hofmann et al. (2023b); Han et al. (2024)) in flat-band systems.
@@ II.1 Quantum metric and coherence length
where $h^{\sigma}_{ij,\alpha\beta}$ is the hopping integral and $U$ denotes the on-site attractive interaction strength. $a_{i\alpha\sigma}$ annihilates a fermion with spin $\sigma$ in the orbital $\alpha$ at the site $i$ (we may call $a_{i\alpha\sigma}$ orbital fermions). Considering an isolated band near the Fermi energy separated from other bands with a large band gap, we can have an effective one-band description. For $s$-wave superconducting phase, it is common to introduce orbital-dependent order parameters $\Delta_{\alpha}=-U\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}\rangle$. The mean-field ground state has been extensively investigated, particularly with regard to the superfluid weight determined by the quantum metric Peotta and Törmä (2015); Törmä et al. (2022). It is possible to project the orbital fermion $a_{i\alpha\sigma}$ onto the fermion $c_{\sigma}$ of the isolated band, which is referred as the band fermion. In particular, we employ the following projection scheme
@@
with $\Delta=1/N\sum_{\alpha\bm{k}}\Delta_{\alpha}|u_{\alpha}(\bm{k})|^{2}$. The projected mean-field Hamiltonian $H_{\mathrm{mf}}$ is independent of the choice of orbital positions $\{\bm{\delta}_{\alpha}\}$. To facilitate the theoretical analysis, we can adopt the uniform pairing condition and the minimal quantum metric Huhtinen et al. (2022b). The former assumes that the pairing potentials are the same for different orbitals, and the latter is specific to orbital positions corresponding to the minimal trace of quantum metric. Then we can define the Cooper pair operator $\hat{\Delta}(\bm{q})=\frac{1}{N}\sum_{i\alpha}e^{-i\bm{q}\cdot(\bm{r}_{i}+\bm{\delta}_{\alpha})}a_{i\alpha\downarrow}a_{i\alpha\uparrow}$ which is formulated after projection as
@@
Here the form factor $\Lambda(\bm{k}+\bm{q},\bm{k})=\sum_{\alpha}u^{*}_{\alpha}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})$ appears as the overlap between two Bloch states. Then we can evaluate the pairing correlator $\mathcal{C}(\bm{r})=\sum_{\bm{q}}e^{-i\bm{q}\cdot\bm{r}}\langle\hat{\Delta}(\bm{q})\hat{\Delta}^{\dagger}(\bm{q})\rangle$ to deduce the coherence length. The pairing correlator $\mathcal{C}(\bm{r})$ is expected to decay exponentially not as a function of $|\bm{r}|$ at zero temperature for an isotropic system. In other words, $\mathcal{C}(\bm{r})\sim e^{-|\bm{r}|/\xi}$ and the decay length $\xi$ is the superconducting coherence length Annett (2004). As shown in Supplementary Note II, $\mathcal{C}(\bm{r})\equiv\sum_{\bm{q}}e^{-i\bm{q}\cdot\bm{r}}\mathcal{M}(\bm{q})$, where
Here, $G(i\omega_{n},\bm{k})$ is the Gor’kov’s Green function of the band fermions $c_{\sigma}$ and $\omega_{n}=(2n+1)\pi T$ is the Matsubara frequency, as defined in Methods section. Then we can extract the coherence length by $\xi^{2}=-\frac{1}{2\mathcal{M}(0)}\left.\frac{d^{2}\mathcal{M}(\bm{q})}{dq^{2}}\right|_{q=0}$ with $q=|\bm{q}|$ at zero temperature. It is essential to emphasize that the validity of the expression in Eq. (8) hinges on the uniform pairing condition, specifically in relation to the Bloch states of the minimal quantum metric Huhtinen et al. (2022b). Without these conditions, the coherence length calculated using Eq. (8) will be overestimated, and additional details can be found in the Supplementary Note II. To see how the quantum metric affects the coherence length $\xi$, the form factor $\Lambda$ enters $\mathcal{M}(\bm{q})$ such that
@@
Figure 3: a The energy spectrum of the flat-band model in Ref. Hofmann et al. (2023b). The solid (dashed) lines denote $t_{2}=0$ and $t_{2}=0.02t$, respectively. $t_{2}$ denotes the nearest hopping which makes the band dispersive. b The profiles of the quantum metric $\mathrm{Tr}[\mathcal{G}]$ of the conduction band in the first Brillouin zone. c The calculated coherence length $\xi$ for $\chi=5$ as a function of the attractive interaction $U/t$. The red, purple and blue denote the cases of $t_{2}=0$, $t_{2}=0.01t$ and $t_{2}=0.02t$, respectively. The theoretical bound $\ell_{\mathrm{qm}}$ is indicated by a dashed green line which coincides with the red one. d The quantum metric dependence of $\xi$ as the parameter $\chi$ varies when $U=0.4t$. The dashed light blue line marks the length scale $\ell_{\mathrm{qm}}$. All calculations are conducted at $k_{B}T=0.001t$ and at half-filling $\mu=t$.

### v6

@@ Abstract
The coherence length $\xi$ is the fundamental length scale of superconductors which governs the sizes of Cooper pairs, vortices, Andreev bound states, and more. In BCS theory, the coherence length is $\xi_{\mathrm{BCS}}=\hbar v_{F}/\Delta$, where $v_{F}$ is the Fermi velocity and $\Delta$ is the pairing gap. It is clear that increasing $\Delta$ will shorten $\xi_{\mathrm{BCS}}$. In this work, we show that the quantum metric, which is the real part of the quantum geometric tensor, gives rise to an anomalous contribution to the coherence length. Specifically, $\xi=\sqrt{\xi_{\mathrm{BCS}}^{2}+\ell_{\mathrm{qm}}^{2}}$ for a superconductor where $\ell_{\mathrm{qm}}$ is the quantum metric contribution. In the flat-band limit, $\xi$ does not vanish but is bound below by $\ell_{\mathrm{qm}}$. We demonstrate that under the uniform pairing condition, $\ell_{\mathrm{qm}}$ is controlled by the quantum metric of minimal trace in the flat-band limit. Physically, the Cooper pair size of a superconductor cannot be squeezed down to a size smaller than $\ell_{\mathrm{qm}}$ which is a fundamental length scale determined by the quantum geometry of the wave functions. Lastly, we compute the quantum metric contributions for the family of superconducting moiré graphene materials, demonstrating the significant role played by quantum metric effects in these narrow-band superconductors.
@@ I Introduction
Bardeen-Cooper-Schriffer (BCS) theory Bardeen et al. (1957) of superconductivity stands as one of the most successful and influential theories in modern physics. It offers a mean-field, yet non-perturbative and microscopic framework for understanding superconductivity. It has been very successful in describing a large number of superconductors Carbotte (1990); Sigrist and Ueda (1991); Blatter et al. (1994). Deviations from the BCS theory are not unusual, which are often attributed to strong interaction effects Stewart (1984); Keimer et al. (2015). Recently, the observations of superconductivity in twisted bilayer graphene Cao et al. (2018); Yankowitz et al. (2019); Arora et al. (2020); Oh et al. (2021); Tian et al. (2023) and related graphene family Liu et al. (2020); Park et al. (2021, 2022) hinted that a new theory is needed to describe superconductors with nearly flat bands. It was observed in a recent experiment Tian et al. (2023) that some important physical quantities deviate greatly from BCS predictions and the microscopic origins behind them are not yet clear.
For example, the BCS superconducting coherence length $\xi_{\mathrm{BCS}}$ is expressed as $\hbar v_{F}/\Delta$, where $v_{F}$ is the Fermi velocity and $\Delta$ is the pairing gap. When the moiré band of twisted bilayer graphene is nearly flat with $v_{F}\approx 10^{3}\text{m}/\text{s}$ and $\Delta\approx 0.2$ meV, $\xi_{\mathrm{BCS}}$ is estimated to be around $3$ nm which is more than one order of magnitude shorter than the values measured using upper critical field measurements Tian et al. (2023). Furthermore, the low Fermi velocity (or equivalently, large effective mass) should lead to a low superfluid stiffness. This results in an expected Berezinskii-Kosterlitz-Thouless transition temperature much lower than the transition temperature measured at optimal doping Tian et al. (2023). It had been pointed out by previous works that the quantum metric Hu et al. (2019); Julku et al. (2020) of the flat bands, which is the real part of the quantum geometric tensor, is essential in sustaining a supercurrent Peotta and Törmä (2015); Törmä et al. (2022). Apart from the investigations of superfluid weight Julku et al. (2016); Liang et al. (2017a, b); Iskin (2018a, b, c); Mondaini et al. (2018); Iskin (2019); Xie et al. (2020); Verma et al. (2021); Herzog-Arbeitman et al. (2022a); Kitamura et al. (2022); Huhtinen et al. (2022); Mao and Chowdhury (2024); Hofmann et al. (2023); Mao and Chowdhury (2023), the quantum geometry affects other physical quantities such as the intrinsic nonlinear transport Gao et al. (2023); Wang et al. (2023); Kaplan et al. (2024) and electron-phonon coupling Yu et al. (2024).
@@
Explicitly, considering the Bloch states of a band represented by $|u(\bm{k})\rangle$, the quantum geometric tensor $\mathfrak{G}$  Provost and Vallee (1980); Berry (1984) is
@@
Here, $a$ and $b$ represent the momentum directions. The quantum geometric tensor can be decomposed into real and imaginary parts as $\mathfrak{G}=\mathcal{G}-i\mathcal{F}/2$, where the real part $\mathcal{G}$ is the quantum metric and the imaginary part $\mathcal{F}$ is the Berry curvature. Berry curvature arises from the phase difference between adjacent Bloch states and characterizes the band topology of materials Klitzing et al. (1980); Thouless et al. (1982); Bellissard et al. (1994); Hasan and Kane (2010); Qi and Zhang (2011); Shapere and Wilczek (1989). The study of the physical consequences of the Berry curvature has been one of the central topics in modern physics. On the other hand, the effect of quantum metric, which measures the distance between two quantum states Anandan and Aharonov (1990), is much less studied. It was pointed out that the quantum metric provides the size (or the so-called quadratic spread) $\ell_{\mathrm{qm}}$ of the optimally localized Wannier state of a band Marzari and Vanderbilt (1997), where $\ell_{\mathrm{qm}}=\sqrt[4]{\det\overline{\mathcal{G}}}$. Here, $\overline{\mathcal{G}}$, defined in Eq. (11), is the weighted average of the quantum metric of the Bloch states within a band. The mathematical definition of $\ell_{\mathrm{qm}}$, which we call the quantum metric length, measures the minimal spread of the Wannier functions, is schematically illustrated in Fig. 1. However, the impact of the quantum metric length $\ell_{\mathrm{qm}}$ on physical quantities was not clear. Until very recently, the Ginzburg-Landau theory Chen and Law (2024) shown that at zero temperature, $\xi=\ell_{\mathrm{qm}}$ for an exactly flat band, which is independent of the interaction strength.
@@
In other words, there is an anomalous quantum metric contribution to the superconducting coherence length (recall that $\xi_{\mathrm{BCS}}=\hbar v_{F}/\Delta$). In the flat-band limit with vanishing $v_{F}$, the quantum metric effect can be significant and even dominant. We show that this is indeed the case for several moiré superconductors with nearly flat bands Tian et al. (2023); Liu et al. (2020); Park et al. (2021, 2022). Our result gives a possible explanation for why the observed superconducting coherence length in the recent experiment Tian et al. (2023) is much larger than expected. It is worth noting that the coherence length is lattice-geometry independent while the quantum metric is lattice-geometry dependent Simon and Rudner (2020); Huhtinen et al. (2022). To resolve the discrepancy, we apply the uniform pairing condition when evaluating the pair correlators and then demonstrate that $\ell_{\mathrm{qm}}$ is related to the quantum metric of the minimal trace Huhtinen et al. (2022). We delineate the physical picture that, in the presence of the quantum metric, increasing the attractive interaction strength between electrons can only reduce the BCS part of the coherence length and squeeze the Cooper pair size down to the quantum metric length $\ell_{\mathrm{qm}}$, but not further, as demonstrated in Fig. 2.
Figure 2: Bound of the coherence length by quantum metric. a For a conventional superconductor with a dispersive band (as illustrated by the insert) without quantum metric, the coherence length $\xi=\hbar v_{F}/\Delta$ decreases as $\Delta$ ($\Delta$ is the superconducting pairing gap) increases and $\xi$ is not bounded from below. b In the presence of quantum metric, the superconducting coherence length $\xi$ has a lower bound of $\ell_{\mathrm{qm}}$. For a superconductor with a narrow band (as illustrated in the insert), the conventional contribution can be suppressed as $\Delta$ increases.
Additionally, for a topological flat band with nontrivial (spin) Chern number, $\xi\geq a\sqrt{|C|/4\pi}$, where $a$ is the lattice constant as illustrated in Fig. 4. At the end of this work, we show that the quantum metric length $\ell_{\mathrm{qm}}$ is important for the superconducting moiré graphene family. As the new length scale $\ell_{\mathrm{qm}}$ defined by the quantum metric is a fundamental property of the band structure and its importance should be manifested beyond superconducting phenomena, we expect that $\ell_{\mathrm{qm}}$ also plays a crucial role in other interaction-driven ordered states (such as the magnetic or density-wave states Hofmann et al. (2023); Han et al. (2024)) in flat-band systems.
@@ II.1 Quantum metric and coherence length
where $h^{\sigma}_{ij,\alpha\beta}$ is the hopping integral and $U$ denotes the on-site attractive interaction strength. $a_{i\alpha\sigma}$ annihilates a fermion with spin $\sigma$ in the orbital $\alpha$ at the site $i$ (we may call $a_{i\alpha\sigma}$ orbital fermions). Considering an isolated band near the Fermi energy separated from other bands with a large band gap, we can have an effective one-band description. For $s$-wave superconducting phase, it is common to introduce orbital-dependent order parameters $\Delta_{\alpha}=-U\langle a_{i\alpha\downarrow}a_{i\alpha\uparrow}\rangle$. The mean-field ground state has been extensively investigated, particularly with regard to the superfluid weight determined by the quantum metric Peotta and Törmä (2015). It is possible to project the orbital fermion $a_{i\alpha\sigma}$ onto the fermion $c_{\sigma}$ of the isolated band, which is referred as the band fermion. In particular, we employ the following projection scheme
@@
with $\Delta=1/N\sum_{\alpha\bm{k}}\Delta_{\alpha}|u_{\alpha}(\bm{k})|^{2}$. The projected mean-field Hamiltonian $H_{\mathrm{mf}}$ is independent of the choice of orbital positions $\{\bm{\delta}_{\alpha}\}$. To facilitate the theoretical analysis, we can adopt the uniform pairing condition and the minimal quantum metric Huhtinen et al. (2022). The former assumes that the pairing potentials are the same for different orbitals, and the latter is specific to orbital positions corresponding to the minimal trace of quantum metric. Then we can define the Cooper pair operator $\hat{\Delta}(\bm{q})=\frac{1}{N}\sum_{i\alpha}e^{-i\bm{q}\cdot(\bm{r}_{i}+\bm{\delta}_{\alpha})}a_{i\alpha\downarrow}a_{i\alpha\uparrow}$ which is formulated after projection as
@@
Here the form factor $\Lambda(\bm{k}+\bm{q},\bm{k})=\sum_{\alpha}u^{*}_{\alpha}(\bm{k}+\bm{q})u_{\alpha}(\bm{k})$ appears as the overlap between two Bloch states. Then we can evaluate the pairing correlator $\mathcal{C}(\bm{r})=\sum_{\bm{q}}e^{-i\bm{q}\cdot\bm{r}}\langle\hat{\Delta}(\bm{q})\hat{\Delta}^{\dagger}(\bm{q})\rangle$ to deduce the coherence length. The pairing correlator $\mathcal{C}(\bm{r})$ is expected to decay exponentially as a function of $|\bm{r}|$ at zero temperature for an isotropic system. In other words, $\mathcal{C}(\bm{r})\sim e^{-|\bm{r}|/\xi}$ and the decay length $\xi$ is the superconducting coherence length Annett (2004). As shown in Supplementary Note 2, $\mathcal{C}(\bm{r})\equiv\sum_{\bm{q}}e^{-i\bm{q}\cdot\bm{r}}\mathcal{M}(\bm{q})$, where
$\!\mathcal{M}(\bm{q})\!=\!\frac{T}{N}\!\sum_{n\bm{k}}\!|\Lambda(\bm{k}+\bm{q},\bm{k})|^{2}G_{0}(i\omega_{n},\bm{k}+\bm{q})G_{0}(-i\omega_{n},-\bm{k}).$ (8)
Here, $G_{0}(i\omega_{n},\bm{k})$ is the normal Gor’kov’s Green function of the band fermions $c_{\sigma}$ and $\omega_{n}=(2n+1)\pi T$ is the Matsubara frequency, as defined in Methods section. Then we can extract the coherence length by $\xi^{2}=-\frac{1}{2\mathcal{M}(0)}\left.\frac{d^{2}\mathcal{M}(\bm{q})}{dq^{2}}\right|_{q=0}$ with $q=|\bm{q}|$ at zero temperature. It is essential to emphasize that the validity of the expression in Eq. (8) hinges on the uniform pairing condition, specifically in relation to the Bloch states of the minimal quantum metric Huhtinen et al. (2022). Without these conditions, the coherence length calculated using Eq. (8) will be overestimated, and additional details can be found in the Supplementary Note 2. To see how the quantum metric affects the coherence length $\xi$, the form factor $\Lambda$ enters $\mathcal{M}(\bm{q})$ such that
@@
Figure 3: Quantum metric and coherence length for topological trivial flat-band model. a The energy spectrum of the flat-band model in Ref. Hofmann et al. (2023). The solid (dashed) lines denote $t_{2}=0$ and $t_{2}=0.02t$, respectively. $t_{2}$ denotes the nearest hopping which makes the band dispersive. b The profiles of the quantum metric $\mathrm{Tr}[\mathcal{G}]$ of the conduction band in the first Brillouin zone. The color bar denotes the magnitude of $\mathrm{Tr}[\mathcal{G}]$. c The calculated coherence length $\xi$ for $\chi=5$ as a function of the attractive interaction $U/t$. The red, purple and blue denote the cases of $t_{2}=0$, $t_{2}=0.01t$ and $t_{2}=0.02t$, respectively. The theoretical bound $\ell_{\mathrm{qm}}$ is indicated by a dashed green line which coincides with the red one. d The quantum metric dependence of $\xi$ as the parameter $\chi$ varies when $U=0.4t$. The dashed light blue line marks the length scale $\ell_{\mathrm{qm}}$. All calculations are conducted at $k_{B}T=0.001t$ and at half-filling $\mu=t$.